A 0.45 kg object connected to a light spring with a force constant of 25.0 N/m oscillates on a frictionless horizontal surface. If the spring is compressed 4.0 cm and released from rest.

(a) Determine the maximum speed of the object. cm/s

(b) Determine the speed of the object when the spring is compressed 1.5 cm. cm/s

(c) Determine the speed of the object when the spring is stretched 1.5 cm. cm/s

(d) For what value of x does the speed equal one-half the maximum speed? cm
Answer #1

angular velocity

w^2 = k/m = 25 / 0.45

w = 7.454 rad/s

==========

a)

using conservation of energy

0.5 m v^2 = 0.5 kx^2

0.45* v^2 = 25* 0.04^2

v = 29.81 cm/s

==========

b)

v^2 = w^2 ( A^2 - x^2)

v^2 = 7.454^2* ( 0.04^2 - 0.015^2)

v = 27.98 m/s

=========

c)

v^2 = 7.454^2* ( 0.04^2 - 0.015^2)

v = 27.98 m/s

=========

d)

v^2 = w^2 ( A^2 - x^2)

(0.2981/2)^2 = 7.454^2* ( 0.04^2 - x^2)

x = 3.464 cm